5/8 of the vehicals in the parking lot are cars. The rest are vans and trucks. If there are 780 cars in the lot how many total vechials are in the lot.
I put 5/8 and x/780 the crossed multiplied and solved from there and got 488 i know that's wrong unless i was suppose add that to the 780. Thank you for your help

Answer by Alan3354(69443)

5/8 of the vehicals in the parking lot are cars. The rest are vans and trucks. If there are 780 cars in the lot how many total vechials are in the lot.
It's vehicles.
===================
I put 5/8 and x/780 the crossed multiplied and solved from there and got 488 i know that's wrong unless i was suppose add that to the 780.
---------------
780 = (5/8)*v
Mutliply by 8
6240 = 5v
v = 1248 vehicles
------------------------
You can do
5/8 = 780/v --> v = 1248